OpenStudy (anonymous):

cos^3θ. sin 3θ + sin^3θ cos 3θ =

OpenStudy (anonymous):

cos^3θ. sin 3θ + sin^3θ cos 3θ =0 cosθ(cos^2θ*sin3θ + sin^3θ)=0 cos^2θsin3θ+sin3θ=0 (1-sin^2θ)sin3θ +sin3θ=0 sin3θ(-sin^2θ+1)=0 1-sin^2θ=0 sin^2θ=1 sinθ= \[\sqrt{1}\] therefore θ=90

OpenStudy (ajprincess):

sin3θ=3sinθ-4sin^3θ cos3θ=4cos^3θ-3cosθ cos^3θ. sin 3θ + sin^3θ cos 3θ =cos^3θ(3sinθ-4sin^3θ)+sin^3θ(4cos^3θ-3cosθ) =3cos^3θsinθ-4sin^3θcos^3θ+4cos^3θsin^3θ-3cosθsin^3θ =3cos^3θsinθ-3cosθsin^3θ =3cosθsinθ(cos^2θ-sin^2θ) =3cosθsinθcos2θ =3/2*2cosθsinθcos2θ =3/2sin2θcos2θ